OpenStudy (anonymous):

ln(x)/ln(1.5) - ln(4)/ln(1.5) = 0.5 solve for x

zepdrix (zepdrix):

\[\Large \frac{\ln x}{\ln1.5}-\frac{\ln4}{\ln1.5}\quad=\quad 0.5\]Hmm I guess we could start by multiplying both sides by ln1.5.\[\Large \ln x- \ln 4 \quad=\quad 0.5\cdot \ln1.5\]Ok with that step?

OpenStudy (anonymous):

Yes

zepdrix (zepdrix):

So next we'll need to apply a few rules of logs, this might seem a little tricky so let's do them one at a time. On the left side, we'll want to apply this rule:\[\Large \color{teal}{\log(a)-\log(b)\quad=\quad \log\left(\frac{a}{b}\right)}\]

zepdrix (zepdrix):

Applying this rule, what will that give us on the left side of our equation? :)

OpenStudy (anonymous):

Log (x)-log (4)

zepdrix (zepdrix):

No no :) `log(x)-log(4)` and `ln(x)-ln(4)` are the same thing in this case. Ignore that part of it, what does it change to (the right side) when we apply the rule? Maybe I should have written the rule like this:\[\Large \color{teal}{\ln a-\ln b\quad=\quad \ln\left(\frac{a}{b}\right)}\]

OpenStudy (anonymous):

I guess I dont quite get what to put in for a and b

zepdrix (zepdrix):

It's just a rule to follow D: So in our case we have a=x, b=4 So it will simplify to,\[\Large \ln x-\ln 4 \quad=\quad \ln\left(\frac{x}{4}\right)\]See how that follows the format of the rule in blue? :o

OpenStudy (anonymous):

Oh ok

zepdrix (zepdrix):

So here's what we've got so far:\[\Large \ln\left(\frac{x}{4}\right) \quad=\quad 0.5\cdot \ln1.5\]

OpenStudy (anonymous):

That we do :)

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

I guess I dont quite get what to put in for a and b

OpenStudy (anonymous):

Log (x)-log (4)

OpenStudy (anonymous):

Yes

zepdrix (zepdrix):

So our next rule we want to apply:\[\LARGE \color{royalblue}{b\cdot \ln(a)\quad=\quad \ln(a^b)}\]We're going to apply this to the right side of our equation. Uh oh the chat is gettin all jumbled up :( weird..

OpenStudy (anonymous):

Why is that the next rule?

zepdrix (zepdrix):

The step we apply `after` this one will get rid of the natural logs, but not if we have a coefficient in front of the log. That 0.5 is causing a problem, so we need a way to bring it inside of the log before we can proceed.

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

Ln(0.5^1.5) ?

zepdrix (zepdrix):

Woops! You have the numbers backwards there.

OpenStudy (anonymous):

Ohh whoops

OpenStudy (anonymous):

Fixed it

zepdrix (zepdrix):

So we get this on the right side, yes? Ln(1.5^0.5)

zepdrix (zepdrix):

So that brings us to this point:\[\Large \ln\left(\frac{x}{4}\right) \quad=\quad \ln\left(1.5^{0.5}\right)\]

zepdrix (zepdrix):

When we raise something to the 0.5 or 1/2 power, that's the same as taking the square root. So we can clean up this expression a tad by rewriting it as a square root.\[\Large \ln\left(\frac{x}{4}\right) \quad=\quad \ln \sqrt{1.5}\]

zepdrix (zepdrix):

The next step would be to notice that both sides have the same operation being applied to them (the natural log), so the contents of those logs will be equal.\[\Large \frac{x}{4}\quad=\quad \sqrt{1.5}\]

OpenStudy (anonymous):

Ln cancel right? Final answer 4.898979486?

zepdrix (zepdrix):

Mmm yah that sounds right! :)

OpenStudy (anonymous):

Awesome! Thanks so much!